# An Effective Incorporating Heterogeneous Knowledge Curriculum Learning for Sequence Labeling

### v1

@@ 3.2.   Difficulty Metrics
Bayesian uncertainty (BU). Buntine and Weigend (1991) proposed that model uncertainty can be assessed by the Bayesian Neural network. According to Wang et al. (2019), the higher predicted probability variance reflects the higher uncertainty that the model has with respect to the sample. Thus, if the model has a large variance in the predicted probability of the sample, the sample is difficult, and vice versa. In this work, we employ the widely used Monte Carlo dropout Gal and Ghahramani (2016) to approximate Bayesian inference.
First, we use Monte Carlo dropout Gal and Ghahramani (2016) to obtain samples of token-level tagging probabilities. We perform $K$ times prediction for each sample in the training set to obtain ${P(y_{i}|x_{i})}_{1}$,…, ${P(y_{i}|x_{i})}_{K}$. Then the expectation of token-level tagging probability can be approximated by
$\mathbb{E}[P(y_{i}|x_{i})]\approx\frac{1}{K}\sum_{k=1}^{K}{P(y_{i}|x_{i})}_{k}$
$var(x_{i},\theta)\approx$ (6)
[…]
$var(\theta)_{aver.}=\frac{1}{M}\sum_{i=1}^{M}var(x_{i},\theta)$
$var(\theta)_{max}=max_{i\in[0,M]}var(x_{i},\theta)$ (8)
$S(\theta)^{BU}=var(\theta)_{max}+var(\theta)_{aver.}$ (9)
3.3.   Training Scheduler
The training scheduler controls the pace of CL. In our work, we use the Root function as the control function, which allocates ample time for the model to learn newly added examples while gradually reducing the number of newly added examples during the training process.
$\lambda\ =\ min(1,\sqrt{\frac{1-\lambda_{0}^{2}}{E_{grow}}\cdot E_{s}+\lambda_{0}^{2}})$ (10)
where $E_{grow}$ represents the number of epochs at which $\lambda$ first reaches 1. Here, $\lambda_{0}>0$ denotes the initial proportion of the easiest training samples, and $E_{s}$ represents the number of training epochs. When $\lambda$ reaches 1, it indicates the model can use the entire training set.
4.   Experiments
4.1.   Dataset and Experimental Configurations
Chinese word segmentation (CWS) and part-of-speech (POS) tagging are representative sequence labeling tasks. So we evaluate our approach using six CWS and POS tagging datasets, including Chinese Penn Treebank version 5.0111https://catalog.ldc.upenn.edu/LDC2005T01, 6.0222https://catalog.ldc.upenn.edu/LDC2007T36, and 9.0333https://catalog.ldc.upenn.edu/LDC2016T13, as well as the Chinese portion of Universal Dependencies(UD)  [^4] https://universaldependencies.org/ and PKU  Jin and Chen (2008). Since the data in the UD dataset is in traditional Chinese, we converted it to simplified Chinese. Regarding the CTB datasets, we follow the same approach as previous works Shao et al. (2017); Tian et al. (2020a) by splitting the data into train/dev/test sets. In the case of PKU, We randomly select 10% of the training data to create the development set. For the UD dataset, we used the official splits of the train/dev/test sets. It is important to note that the UD dataset has two versions: UD1, which consists of 15 POS tags, and UD2, which includes 42 POS tags. The details of the six datasets are given in Table 1.
Datasets CTB5 CTB6 CTB9 UD1 UD2 PKU
Train Sent. 18k 23K 106k 4k 4k 17k
Word 494k 641 1696k 99k 99k 482k
Dev Sent. 350 2K 10k 500 500 1.9k
Word 7k 60K 136k 13k 13k 53k
Test Sent. 348 3K 16k 500 500 3.6k
Word 8k 82K 242k 12k 12k 97k
Table 1: Detail of the six datasets.
Various methods have shown significant progress in fine-tuning Pretrained Language Models (PLM) for sequence labeling tasks Tian et al. (2020a, b); Liu et al. (2021). In this study, we use BERT-based Devlin et al. (2019) and RoBERTa-based Liu et al. (2019) models as encoders, following their default settings. According to Qiu et al. (2020), there is little difference in the decoding effects between Softmax and Conditional Random Fields (CRF) for sequence labeling tasks. Consequently, we opt for Softmax as the decoder. The framework of the basic transfer teacher is BERT+Softmax. As for the student model, we choose two simpler models (BERT/RoBERTa+Softmax) and two representative complex models introduced by  Tian et al. (2020a) and  Tang et al. (2023). In their work,  Tian et al. (2020a) employed a two-way attention mechanism, known as MCATT, to integrate dependency relations, POS labels, and syntactic constituents for the joint CWS and POS tagging task. They employed BERT as the encoder for this setup. On the other hand,  Tang et al. (2023) incorporated syntax and semantic knowledge into sequence labeling tasks using a GCN module called SynSemGCN. They applied RoBERTa as the sequence encoder in their experiment.
The values of $\lambda_{0}$ and $E_{grow}$ in Eq. 10 are set to 0.3 and 10, respectively. For the TLC difficulty metric, the key hyperparameter $N$ is set to 5. For the BU difficulty metric, the key hyperparameter $K$ is set to 3. The number of training epochs for the teacher model denoted as $E_{0}$, is set to 5. For further details on the important hyper-parameters of the model, please refer to Table 2.
Curriculum Learning Baselines. There are two baseline difficulty metrics for curriculum Learning: a. Random, which assigns samples in a random order; b. Sentence length (Length), which ranks samples from shortest to longest based on the intuition that longer sequences are more challenging to encode.
Hyper-parameters
@@ 4.1.   Dataset and Experimental Configurations
  Model TCL Setting CTB5 CTB6 CTB9 UD1 UD2 PKU
CWS POS CWS POS CWS POS CWS POS CWS POS CWS POS
BERT NA 98.49 96.34 97.29 94.71 97.53 94.61 98.20 95.70 98.25 95.33 98.01 95.49
Random 98.77 96.80 97.35 94.83 97.35 94.71 98.26 95.74 98.32 95.51 98.32 96.30
Length 98.88 96.81 97.44 94.88 97.75 95.03 98.34 95.94 98.27 95.35 98.40 96.29
TLC 98.71 96.85 97.41 94.91 97.74 94.99 98.30 95.81 98.25 95.47 98.40 96.33
MNLP 98.90 96.99 97.40 94.92 97.74 95.01 98.27 95.78 98.30 95.58 98.40 96.31
BU 98.84 96.80 97.45 94.92 97.75 95.00 98.28 95.84 98.32 95.60 98.41 96.32
RoBERTa NA 98.66 96.50 97.31 94.80 97.58 94.70 98.25 95.80 98.17 95.32 98.30 95.85
Random 98.70 96.75 97.48 94.90 97.33 94.23 98.34 95.82 98.23 95.42 98.51 96.51
Length 98.73 96.73 97.42 94.91 97.77 95.09 98.26 95.80 98.25 95.47 98.53 96.47
TLC 98.79 96.86 97.41 94.98 97.75 95.03 98.32 95.91 98.20 95.52 98.63 96.50
MNLP 98.72 96.85 97.47 95.06 97.74 95.07 98.23 95.78 98.30 95.50 98.51 96.54
BU 98.77 96.80 97.42 95.00 97.77 95.09 98.21 95.85 98.29 95.60 98.54 96.50
MCATT NA 97.72 96.55 97.39 94.80 97.28 93.92 98.28 95.70 98.28 95.40 98.20 96.25
Random 98.81 96.84 97.37 94.90 97.13 93.83 98.25 95.61 98.31 95.47 98.38 96.27
Length 98.83 96.85 97.35 94.82 97.60 94.71 98.23 95.66 98.34 95.61 98.40 96.25
TLC 98.83 96.89 97.37 94.83 97.60 94.68 98.33 95.61 98.15 95.53 98.41 96.30
MNLP 98.85 96.81 97.41 94.92 97.64 94.70 98.42 95.87 98.26 95.35 98.41 96.30
BU 98.91 96.87 97.42 94.90 97.69 94.88 98.34 95.66 98.18 95.49 98.43 96.32
SynSemGCN NA 98.75 96.73 97.89 94.95 97.70 94.91 98.24 95.93 98.21 95.51 98.37 96.17
Random 98.84 97.86 97.99 95.05 97.50 94.54 98.25 95.85 98.22 95.44 98.48 96.40
Length 98.80 96.84 97.40 94.94 98.01 95.54 98.22 95.89 98.24 95.57 98.53 96.48
TLC 98.83 97.81 97.98 95.02 98.00 95.53 98.36 95.99 98.28 95.50 98.61 96.55
MNLP 98.78 97.72 98.04 95.13 98.02 95.53 98.26 95.89 98.30 95.61 98.56 96.48
BU 98.90 97.95 98.05 95.14 98.03 95.48 98.34 96.06 98.23 95.48 98.59 96.54
Table 3: Experimental results of different models using different TCL settings on test sets of six datasets. SynSemGCN represents the model from Tang et al. (2023) and MCATT denotes the model from Tian et al. (2020b). Here, “CWS” represents the F1 value of CWS, and “POS” means the F1 value of the joint CWS and POS tagging. “NA” means without the TCL training strategy. The maximum F1 scores for each dataset are highlighted.
  Model CTB5 CTB6 CTB9 UD1 UD2 PKU
CWS POS CWS POS CWS POS CWS POS CWS POS CWS POS
Chen et al. (2017) - 93.19 - - - - - - - - - 90.16
Shao et al. (2017) 98.02 94.38 - - 96.67 92.34 95.16 89.75 95.09 89.42 - -
Zhang et al. (2018) 98.50 94.95 96.36 92.51 - - - - - - 96.35 94.14
Zhao et al. (2020) 99.04 96.11 97.75 94.28 97.75 94.28 - - - - 97.48 95.35
Tian et al. (2020b)(MCATT) 98.73 96.60 97.30 94.74 97.69 94.78 98.29 95.50 98.27 95.38 - -
Tian et al. (2020a)(TWASP) 98.77 96.77 97.43 94.82 97.75 94.87 98.32 95.60 98.33 95.46 - -
Liu et al. (2021)(LEBERT) - 97.14 - - - - - 96.06 - 95.74 - -
Tang et al. (2023)(SynSemGCN) 98.83 96.77 97.86 94.98 97.67 94.92 98.31 95.83 98.25 95.50 98.05 95.50
BERT+TCL(BU) 98.84 96.80 97.45 94.92 97.75 95.00 98.28 95.84 98.32 95.60 98.41 96.32
RoBERTa+TCL(BU) 98.77 96.80 97.42 95.00 97.77 95.09 98.21 95.85 98.29 95.60 98.54 96.50
MCATT+TCL(BU) 98.91 96.87 97.42 94.90 97.69 94.88 98.34 95.66 98.18 95.49 98.43 96.32
SynSemGCN+TCL(BU) 98.90 97.95 98.05 95.14 98.03 95.48 98.34 96.06 98.23 95.48 98.59 96.54
Table 4: Performance on the test set of six datasets from previous studies, and different models trained with our approach (using the TCL strategy with the BU difficulty metric).
4.2.   Overall Experimental Results
Table 3 presents the experimental results of four models with different CL settings. The reported F1 values for CWS and the joint task represent the averages obtained from three experiments. The experimental results reveal several noteworthy conclusions.
Firstly, the TCL methodology proposed in this paper exhibits flexibility as it can be integrated with both simple and complex models. It is evident that the TCL approach enhances the performance of all four models in terms of F1 values, which indicates the effectiveness of our approach.
Secondly, we compare our approach with various difficulty metrics in the TCL strategy, including Random, Length, TLC, MNLP, and BU. As observed in Table 3, the difficulty metrics proposed in this study outperform Random and Length metrics across most datasets. Moreover, the BU metric performs the best on the majority of datasets, outperforming models utilizing TLC and MNLP metrics. In particular, the SynSemGCN model, utilizing the TCL approach with the BU difficulty metric, achieves a 1.22% improvement in POS tagging F1 value on the CTB5 dataset compared to the model without TCL. Furthermore, it is worth noting that the impact of the TCL strategy is more pronounced when applied to larger datasets compared to smaller ones. Specifically, the datasets UD1 and UD2 exhibit less notable improvement compared to the CTB series datasets. This observation aligns with the finding from a previous study conducted by Wang et al. (2022), which supports the effectiveness of CL in addressing data heterogeneity. Therefore, it is particularly beneficial for datasets like CTB that encompass a wide range of diverse text types.
Finally, we compare our model with previous methods that incorporate external knowledge or resources into the encoder, such as lexicons Liu et al. (2021), orthographical features Shao et al. (2017), n-grams Tian et al. (2020b); Zhao et al. (2020); Zhang et al. (2018); Chen et al. (2017), and syntax knowledge Tian et al. (2020a); Tang et al. (2023). Table 4 lists the comparison of different models, where X+CL(BU) indicates training the model using the proposed TCL method with the BU difficulty metric. The results demonstrate that models using CL exhibit further enhancements in performance, outperforming previous methods across the majority of datasets.
4.3.   Effect of Two-stage Curriculum Learning
  Model CTB5 PKU
CWS POS CWS POS
Ours 98.90 97.95 98.59 96.54
w/o data CL(BU) 98.90 97.88 98.53 96.32
w/o model CL(BU) 98.85 97.51 98.41 96.20
w/o TCL 98.75 96.73 98.37 96.17
Table 5: Ablation experimental results of TCL. The baseline model “w/o CL” denotes the model SynSemGCN; “w/o model CL” means the student model always uses the initial data order sorted by the transfer teacher model; “w/o data CL” indicates the initial training samples for the student model is drawn randomly from the training set; “Ours” indicates “SynSemGCN+CL(BU)”. All models with CL in this table use BU as the difficulty metric.
In this section, we discuss the influence of TCL. We perform ablation studies by removing either the data-level CL or the model-level CL. The results of ablation experiments are summarized in Table 5. Notably, both curricula prove effective within the model and complement each other. Consequently, it is reasonable to incorporate a teacher transfer model that guides the student model during the initial stages of self-paced learning.
In particular, model-level CL contributes more significantly than data-level CL. This observation is intuitive, as model-level CL plays a more significant role throughout the entire training process of the student model, whereas data-level CL is primarily involved in the initial stages of training the student model.
Additionally, we compare the training time of models with and without TCL. The experimental results are presented in Table 6, where all models were trained for 50 epochs, and the time of models using TCL includes the time spent on training the teacher model and calculating the difficulty values for the student model.
Analyzing the results, it becomes evident that TCL reduces training time significantly. This is because only selected data are involved in training during the model-level CL phase. Notably, TCL not only enhances model performance but also reduces training time for both models by over 20%.
  Model TCL(BU) CTB5
CWS POS Time
MCATT × 98.72 96.55 645m
98.91 96.87 507m
SynSemGCN × 98.75 96.73 393m
98.90 97.95 287m
Table 6: Comparison of experimental results and training time (in minutes) on CTB5 dataset for baseline models MCATT and SynSemGCN with and without CL(BU).
4.4.   Comparison of Difficulty Metrics
In this section, we investigate the impact of various difficulty metrics during the model-level CL training process for the SynSemGCN model.
Figure 2 depicts the change in F1 scores on the PKU dataset development set over the initial 10 epochs of the model-level CL process. After 10 epochs, all training data are employed, so the initial 10 epochs offer insights into the influence of different metrics.
Observing the figure, we can discern that the four alternative difficulty metrics outperform the Random metric. Notably, a significant performance gap is evident between the Length metric and the three metrics introduced in this paper. BU, in particular, achieves the highest performance, suggesting that uncertainty-based metrics can select samples that align better with the model’s learning path, resulting in accelerated learning for the model.
4.5.   Effect of Hyper-parameters
@@
In this section, we explore the impact of the hyperparameters on the performance of CL. The adjustment of the parameters is based on the SynSemGCN+TCL(BU) model.
First, we investigate the impact of the hyper-parameter $\lambda_{0}$ on CL performance. We conduct the experiments on the CTB5 dataset, tuning the value of $\lambda_{0}$ in the model-level pacing function Eq. 10, and the experimental results are represented by a line graph as shown in Figure 3. As observed, the model achieves optimal performance when $\lambda_{0}=0.3$. However, when the value exceeds 0.4, the model’s performance gradually deteriorates.
@@
Finally, we explore the impact of different $K$ values on the BU difficulty metric, which determines the number of dropout times. The experiments are conducted on the CTB6 dataset, and the results are summarized in Table 9. Notably, the model achieves optimal performance when $K=3$. Therefore, we select $K=3$ for all the above experiments.
Due to space constraints, we do not report the experimental results for the variation of $N$ in the TLC metric. However, we can state that the best performance was experimentally demonstrated when $N=5$. In the future, we will add these experimental results in the appendix.
  Para. CTB6 PKU
CWS POS CWS POS
$K$=2 98.01 95.11 98.55 96.53
$K$=3 98.05 95.14 98.59 96.54
$K$=4 98.03 95.13 98.52 96.46
Table 9: The effect of of $K$ times dropout in BU difficulty metric.
5.   Conclusion
This paper introduces a novel two-stage curriculum learning framework aimed at enhancing performance and accelerating the training process for sequence labeling tasks. Focusing on the sequence labeling task of CWS and POS tagging, this framework demonstrates its effectiveness in improving results. To further broaden the scope of this framework, future research could explore its application to other sequence labeling tasks, such as named entity recognition
There are some limitations. First, our approach may not yield significant improvement when applied to small datasets. Additionally, the design of our difficulty metrics involves the setting of many hyperparameters, which may pose a challenge in terms of optimization.
6.   Bibliographical References

### v2

@@ Appendix B Difficulty Metric Baselines
Appendix C Parameters Setting
The key experimental parameter settings are shown in Table 7.
Hyper-parameters Value
@@ Appendix C Parameters Setting
Appendix D Effect of Hyper-parameters
@@
  Para. CTB6 PKU
CWS POS CWS POS
$K$=2 98.17 95.43 98.73 96.58
$K$=3 98.10 95.59 98.77 96.97
$K$=4 98.09 95.56 98.69 96.38
Table 10: The effect of of $K$ times dropout in BU difficulty metric.
In this section, we explore the impact of the hyperparameters on the performance of DCL. The adjustment of the parameters is based on the SynSemGCN+DCL(BU) model.
First, we investigate the impact of the hyper-parameter $\lambda_{0}$ on DCL performance. We conduct the experiments on the CTB5 dataset, tuning the value of $\lambda_{0}$ in the model-level pacing function Eq. 6, and the experimental results are represented by a line graph as shown in Figure 3. As observed, the model achieves optimal performance when $\lambda_{0}=0.3$. However, when the value exceeds 0.4, the model’s performance gradually deteriorates.
@@
Then, we explore the impact of different $K$ values on the BU difficulty metric, which determines the number of dropout times. The experiments are conducted on the CTB6 dataset, and the results are summarized in Table 10. Notably, the model achieves optimal performance when $K=3$. Therefore, we select $K=3$ for all the above experiments.
Finally, we evaluate the effect of varying $N$ in the TLC metric. As shown in Table 11, the best performance is achieved when $N=5$.
  Para. PKU
CWS POS
$N$=1 98.55 96.49
$N$=2 98.53 96.49
$N$=3 98.56 96.52
$N$=4 98.55 96.51
$N$=5 98.71 96.64
$N$=6 98.52 96.52
$N$=7 98.56 96.51
$N$=8 98.55 96.48
$N$=9 98.53 96.49
$N$=10 98.55 96.51
Table 11: The impact of $N$ in TLC difficulty metric.
                      Models CTB5
A B CWS POS
BERT+DCL(BU) BERT > >
BERT+DCL(MNLP) BERT > >
BERT+DCL(TLC) BERT > >
BERT+DCL(BU) BERT+CL(Length) > >
BERT+DCL(MNLP) BERT+CL(Length) > >
BERT+DCL(TLC) BERT+CL(Length) > >
Table 12: Statistical significance test of F-score for our method and baselines on the CTB5 dataset.
Appendix E Statistical Significance Test
In this section, we conduct significance testing experiments. Following  Wang et al. (2010), we use the bootstrapping method proposed by  Zhang et al. (2004), which is operated as follows. In this process, starting with a test set $T_{0}$ comprising $N$ test examples, we repeatedly sample $N$ samples from $T_{0}$ to form $T_{1}$ and then repeat the process for $M$ times to form the test set collection, $\{T_{1},T_{2},...,T_{M}\}$, where $M$ is set to 1000 in our testing procedure. Two systems denoted as $A$ and $B$, are assessed on the initial test set $T_{0}$, resulting in scores $a_{0}$ and $b_{0}$, respectively. The disparity between the two systems, labeled as $\delta_{0}$, is calculated as $\delta_{0}=a_{0}-b_{0}$. Repeating this process for each test set produces a set of $M$ discrepancy scores, denoted as $\{\delta_{0},\delta_{1},...,\delta_{M}\}$.
Following the methodology proposed by  Zhang et al. (2004), we compute the 95% confidence interval for the discrepancies (i.e., the 2.5th percentile and the 97.5th percentile) between the two models. If the confidence interval does not overlap with zero, it is affirmed that the differences between systems A and B are statistically significant  Zhang et al. (2004).